Inventory 30 POINTS Compute the inventory methods for TV Vison. Suppose TV Vison started March with an inventory of 50 plasma TVs that cost $2,010 each, for a total beginning inventory value of $100,500.

During March, the firm made the following purchases:

March 2 200 TVs for $2,000 each

March 10 150 TVs for $1,800 each

March 20 100 TVs for $1,500 each

March 29 50 TVs for $1,000 each

During March, the firm made the following sales:

March 5 110 TVs for $4,000 each

March 12 160 TVs for $4,000 each

March 25 150 TVs for $4,000 each

Instructions:

A] Using periodic inventory record keeping, calculate the cost of goods sold for the month and the ending inventory at the end of the month. Do these calculations using THREE METHODS, Weighted Average Cost, FIFO, and LIFO.

B] All other operating expenses amount to $250,000. Calculate net income using each of the three methods.

C] Using the perpetual inventory record keeping, calculate the ending inventory and cost of goods sold for TV Vison adopting the LIFO and FIFO methods.

D] Why do companies avoid the perpetual weighted moving average method of calculating inventory costs?
Answer #1

A) Weighted Average Cost Using Periodic Inventory

Under Weighted average cost method, weighted average cost per unit is used for calculating cost of goods sold and cost of ending inventory.

Weighted Average cost per unit = Total cost of goods available for sale/Total units available for sale

Total Units purchased = 200+150+100+50 = 500

Total Units sold = 110+160+150 = 420

Total Units in Ending Inventory = Units in Beginning Inventory+Units purchased-Units sold

= 50+500-420 = 130  

Total Units available = Units in Beginning Inventory+Total Units purchased

= 50+500 = 550 units

Total cost of goods available for sales = Beginning Inventory+Total Purchases

= $100,500+(200 TVs*$2,000)+(150 TVs*$1,800)+(100 TVs*$1,500)+(50 TVs*$1,000)

= $100,500+$400,000+$270,000+$150,000+$50,000 = $970,500

Weighted Average Cost per unit = $970,500/550 units = $1,764.54545 per unit

Cost of goods sold = 420 units*$1,764.54545 = $741,109

Cost of Ending Inventory = 130 units*$1,764.54545 = $229,391

FIFO Using Periodic Inventory

Under FIFO method, the goods purchased first are sold first.

130 TVs under FIFO would include 50 TVs purchased on Mar 29 and remaining 80 TVs purchased on Mar 20. Therefore cost of goods sold include cost of beginning inventory and cost of purchases made on Mar 2, 10 and 20 TVs purchased on Mar 20.

Cost of goods sold = $100,500+(200 TVs*$2,000)+(150 TVs*$1,800)+(20 TVs*$1,500)

= $100,500+$400,000+$270,000+$30,000 = $800,500

Cost of Ending Inventory = (80 TVs*$1,500)+(50 TVs*$1,000)

= $120,000+$50,000 = $170,000

LIFO Using Periodic Inventory

Under LIFO method, the goods purchased last are sold first.

130 TVs under LIFO would include 50 TVs from beginning inventory and remaining 80 TVs purchased on Mar 2. Therefore cost of goods sold include cost of purchases made on Mar 10, 20, 29 and 120 TVs (200-80 ending inventory) purchased on Mar 2.

Cost of goods sold = (120 TVs*$2,000)+(150 TVs*$1,800)+(100 TVs*$1,500)+(50 TVs*$1,000)

= $240,000+$270,000+$150,000+$50,000 = $710,000

Cost of Ending Inventory = $100,500+(80 TVs*$2,000)

= $100,500+$160,000 = $260,500

B) Sales = 420 TVs*$4,000 = $1,680,000

Calculation of Net Income under Each method (Amounts in $)

Weighted Average Periodic Method FIFO Periodic LIFO Periodic
Sales 1,680,000 1,680,000 1,680,000
Cost of Goods Sold (741,109) (800,500) (710,000)
Gross Profit 938,891 879,500 970,000
Operating Expenses (250,000) (250,000) (250,000)
Net Income 688,891 629,500 720,000

C) FIFO Using Perpetual Inventory

Under Perpetual method, the cost of goods sold is calculated after each sale. The cost of goods sold and cost of ending inventory same under FIFO periodic and FIFO perpetual method.

Cost of goods sold = $800,500

Cost of Ending Inventory = $170,000

LIFO Using Perpetual Inventory

Under this method, units purchased on Mar 29 remain in ending inventory as it is purchased after all sales and remaining inventory would be from beginning inventory and units purchased on Mar 2.

Cost of Ending Inventory = Beginning Inventory+Units purchased on Mar 29+30 units purchased on Mar 2

= $100,500+(50 TVs*$1,000)+(30 TVs*$2,000)

= $100,500+$50,000+$60,000 = $210,500

Cost of goods sold = Cost of goods available for sale - Cost of Ending Inventory

= $970,500-$210,500 = $760,000

D) This method is complicated because a new weighted average cost is calculated at each time of purchase. Therefore companies avoid the perpetual weighted moving average method of calculating inventory costs.